OpenStudy (rebeccapink6398):

@texaschic101 :D Convert the decimal to a percent. 6.70 A. 6.7% B. 67.0% C. 0.670% D. 670% I think its D. :D

OpenStudy (shamil98):

yep.

OpenStudy (texaschic101):

to convert a decimal to a percent, start at the decimal and move two spaces to the right, and then put your percent. 6.70 -- move 2 spaces to the right and put your percent = 670%

OpenStudy (rebeccapink6398):

Which percent is equal to the given fraction? ~ 9/10 I think its 90%

OpenStudy (shamil98):

correct again.

OpenStudy (texaschic101):

way to go :)

OpenStudy (rebeccapink6398):

how about this one? Which percent is equivalent to 7 and 1/4 I think its 7.25%

OpenStudy (texaschic101):

7 1/4 = 29/4 = 7.25 = 725%

OpenStudy (rebeccapink6398):

thank you :D

OpenStudy (rebeccapink6398):

how about this one? its my last question :D In a recent election, seven out of every ten voters favored a proposition to improve the city's parks. Which fraction and equivalent percent compares the number of voters who favored the proposition to all voters?

OpenStudy (texaschic101):

7/10 = 0.7 = 70%

OpenStudy (rebeccapink6398):

THX!!! 100% again! UR AWESOME!!

OpenStudy (texaschic101):

glad to help :)

OpenStudy (anonymous):

im having problems with this question

OpenStudy (anonymous):

Research the highest interest rate (APY—annual percent yield) for 2-year and 5-year CDs. Document the company's name, interest rate, and minimum investment. The minimum investment must be less than or equal to $5,000.